OpenStudy (anonymous):

Solve for x: 2/3 (x − 2) = 4x. A. -2/5 B. -5/2 C. 2/5 D. 5

OpenStudy (vortish):

i think your answer is correct

OpenStudy (anonymous):

what answer? I didn't answer it. lol.

OpenStudy (anonymous):

2/3 (x − 2) = 4x 2/3(x) - 2/3(2) =4x

OpenStudy (anonymous):

2/3(x)-4x=4/3

OpenStudy (anonymous):

\[\frac{ 2x-12x }{ 3} = \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

I'm still lost..

OpenStudy (anonymous):

use distributive property on 2/3(x-2)

OpenStudy (anonymous):

and then combine like terms

OpenStudy (anonymous):

 Can you go from there? Basically, keep the goal to get x by itself in your head when you are asked to solve for x. (: